OpenStudy (anonymous):

x & y intercept. 5x-y=-5 & -3x+2y=7

Nnesha (nnesha):

`x-intercept` is a point where line intersect x-axis when y=0 so substitute `y for 0` to find x-intercept `y-intercept` is a point where line cross y-axis when x=0 so substitute `x for 9` to find y-itercept

OpenStudy (ellaella02):

x=1/5y-1 for the first and x=2/3y+-7/3 for the second